A chemist prepares a solution of barium chlorate (Ba(C103)2) by measuring out 1.5x10^2mol of barium chlorate into a 500.mL volumetric flask and filling the flask to the mark with water. Calculate the concentration in of the chemist's barium chlorate solution. Round your answer to significant digits.


1
Expert's answer
2021-08-16T02:46:09-0400

500. mL = 0.500 L

"Molarity=\frac{n}{V}=\frac{1.5×10^2mol}{0.500L}=3.0×10^2M"

It is not mentioned how many significant figures the answer should be rounded to, but accorging to general rules there should be 2 significant figures in this case.

Answer: 3.0 x 102 M


The number of moles given is abnormally big and probably incorrect. Such solution cannot be physically prepared as described (150 moles of barium chlorate weighs about 46 kg and cannot fit in the flask and cannot be dissolved in 0.5 liters of water).